Previous | Calendar |
April |
2015 |
Next | ||
---|---|---|---|---|---|---|
Sunday | Monday | Tuesday | Wednesday | Thursday | Friday | Saturday |
29JD88 |
30JD89at 7:00 MST, LI-7200 *ghg files stop being archived |
31JD90 |
1JD91 |
2JD92 |
3JD93 |
4JD94Gap1 in MRS data file (starts) |
5JD95 |
6JD96 |
7JD97Site Visit (ES, JK, NM, +): Details ES added a CO2 span cylinder to the CPEC200 (as a check of instrument span) Site visit by snow hydrology class (NM) Gap1 in MRS data file (ends) SpB COMP1 |
8JD98SpB COMP2 |
9JD99Gap2 in MRS data file (starts) |
10JD100SpB COMP3 |
11JD101 |
12JD102 |
13JD103 |
14JD104Site Visit (JK): Details |
15JD105Site Visit (SpB): Details , Photos (29) Span cylinder at 1400 psi. N2 at 800 psi at approx 10 MST, reset/rebooted LI-7200 from 10:28-10:30 MST, adjusted span/zero flow in LI-6262 swapped SD card from cloudCAM / SpB Jeep Overcast w/ some sun breaks (plus light snow) |
16JD106some snow |
17JD107MORE snow |
18JD108 |
19JD109 |
20JD110 |
21JD111 |
22JD112Site Visit (JK): Details |
23JD113 |
24JD114 |
25JD115NCTC logger (/dev/ttyS10) drops off data system? tried ddn/dup, but didn't fix it. |
26JD116 |
27JD117Snow (CSATs not working) |
28JD118at approx 11 MST, the flow rate in 6262 drops; something constricting flow? (it seems to recover) Still snowy (CSATs still struggling) |
29JD119Site Visit (JK): Details SpB COMPS Exam at approx 12:00 MST, LI-7200 *ghg files stop being archived |
30JD120 |
1JD121Site Visit (SpB): Details , Photos (33) Span cylinder at 1310 psi. N2 at 500 psi at approx 11:20 MST, connected aspiration/fan to the two REBS T/RH sensors at approx 12 MST, disconnected power from LI-7500A at 2.5m at approx 17 MST, reset/rebooted LI-7200 at 15:30 MST, started ops76: removed Tir variables from NCTC logger (started nctc_05.csi) swapped SD card from towerCAM SpB Jeep Sunny in morning, then thunder/lightning/snow/rain |
2JD122 |